# ModServe: Scalable and Resource-Efficient Large Multimodal Model Serving

### v2

@@ 4. ModServe Design and Implementation
Overview. Figure 13 shows ModServe’s design. A pool of Image Instances handles image preprocessing and encoding of image-text requests. The resulting image tokens are passed to a pool of Text Instances, which performs LLM prefill and decode operations. Text-only requests bypass the image components and are queued directly at the Text Instances. Two pools are managed by the Image and Text Pool Managers.
@@ 4.1. Offline LMM Profile Generation
When onboarding a new LMM, ModServe generates resource-performance profiles by characterizing the image encoder and LLM backend independently. This profiling runs controlled inference workloads with varying model parallelisms (e.g., TP-2 and TP-8), batch sizes, and load (i.e., image tokens per second for image encoders and prompt tokens per second for LLM backends) to capture per-stage performance characteristics. To efficiently model performance across different load conditions, ModServe profiles a set of representative load levels (up to the maximum throughput) and extrapolates the behavior for intermediate loads. The resulting profiles take load, parallelism, and batch size as inputs to predict key performance metrics, including encoding latency for Image Instances and prefill time and TBT for Text Instances. The Pool Managers use these profiles to guide operational decisions (Section 4.2): (1) pool autoscaling to meet latency SLOs without overprovisioning, (2) model sharding that selects optimal TP degrees, and (3) max batch sizing for each stage.
@@ 4.2. Decoupled Resource Management
The Image Pool Manager maintains a pool of Image Instances, which preprocess images on CPU and encode image workloads on GPU for image-text requests. Meanwhile, the Text Pool Manager manages a pool of Text Instances responsible for the prefill and decode stages of both image-text and text-only requests. Based on model profiles (Section 4.1), each manager independently optimizes pool autoscaling, model sharding, and max batch sizing to minimize costs while meeting performance SLOs.
Before their online operation, the initial number of Image Instances () is determined using the median image QPS multiplied by the median image encoding latency. The number of Text Instances () is set as divided by the median number of images per request, based on historical LMM inference traces. If no history is available, ModServe initially overprovisions resources to ensure reliability.
@@
For Image Instances, image token counts are precomputed based on static mapping from image dimensions (Figure 3). Autoscaling of Text Instances is based on text token load in CroAttn models but total tokens in DecOnly models due to homogeneous self-attention across modalities. Advanced autoscaling hysteresis prevention techniques (Zou et al., 2024) can be employed to avoid excessive scaling actions caused by transient workload fluctuations but are not covered in this paper.
@@
When scaling beyond a single GPU server, ModServe prioritizes autoscaling over pipeline parallelism (PP) (Shazeer et al., 2020) to maximize throughput, seamlessly transitioning to batch-level optimizations as needed.
@@ 4.3. Modality-Aware LMM Request Serving
Instead, ModServe routes requests by input modality. Image-text requests are assigned to Image Instances with the least image-token load. Large requests (i.e., those with more images) are consequently distributed across multiple Image Instances for parallel processing and encoding (Section 3.1), preventing degraded batching performance that would occur if all images were routed to a single instance. This effectively enables a form of request chunking (Agrawal et al., 2024), where images in a large request can be processed in an interleaved manner with other requests, reducing HoL blocking and improving scheduling flexibility.
@@
Pool Managers continuously monitor SLO attainment and trigger pool autoscaling when the rate falls below a predefined threshold (default 0.99 with a sensitivity study in Section 5.3), ensuring adaptive resource allocation under dynamic workloads (especially in cases of unpredictable traffic). ModServe can work with state-of-the-art batch scheduling techniques (Agrawal et al., 2024; Patel et al., 2024) to optimize TBT during the decode stage, which we leave to future work as we do not observe TBT degradation in LMM characterization (Section 3.1).
Image Token Transfer. Once image processing is complete, ModServe transfers image tokens from Image Instances to Text Instances via a pull-based RDMA mechanism. Unlike a push-based approach that immediately sends image tokens upon availability, this strategy reduces synchronization overhead and optimizes load balancing.
Since image-text requests may involve multiple images processed across different Image Instances, the transfer follows a many-to-one pattern, aggregating all image tokens before text processing begins. This defers transfer until all tokens are ready, allowing for better scheduling decisions by routing requests to the least-loaded Text Instance. Meanwhile, queuing at the Text Instance overlaps with token transfer, effectively hiding transfer latency.
@@ 4.4. Implementation
We implement ModServe using 5,000 lines of Python code. We base the Text Instance on vLLM (Kwon et al., 2023) (v0.7.2), a state-of-the-art generative model inference platform, and build the Image Instance on HuggingFace Transformers (HuggingFace, 2024b). The modular architecture of ModServe enables easy integration with other serving engines (e.g., TensorRT (NVIDIA, 2024) and DeepSpeed (Aminabadi et al., 2022)). We use numactl to restrict CPU and memory usage of image preprocessing to a single NUMA node, which reduces memory access latency and performance variation. To ensure efficient GPU-to-GPU memory transfer of image tokens, we use PyTorch’s distributed communication with the NCCL backend and GPU Direct RDMA. The P99 transfer latency of image tokens per request is 5 ms.
@@ 5.1. Experimental Setup
Models and Workloads. We use Llama3.2-11B and InternVL-26B as representative models for CroAttn and DecOnly LMMs, respectively. To ensure realistic workload distribution, we reuse the LMM dataset from Section 3.1 and adopt the inter-arrival timestamps of requests and the number of images associated with each request (ranges from 0 to 16) from the production LMM inference trace (Section 3.2).
@@
SLO Definition. We define the SLO metrics for LMM inference based on the TTFT/TBT during the isolated run of a single text-only request and text-image request (with one image) on the monolith baseline setup. Then, we scale the SLO metrics with a constant factor (i.e., SLO factor) to evaluate how ModServe performs under tight/relaxed SLOs (Section 5.3). The SLOs are defined on P99 tail latency.
@@ 5.2. End-to-end Performance
Figure 16 compares the number of GPUs used by ModServe and vLLM (monolith) to serve the image-driven burst in the production trace. ModServe takes 41.3% and 25% fewer GPUs compared to vLLM to serve Llama-3.2 (CroAttn) and InternVL (DecOnly) models respectively while meeting the tail latency SLOs. ModServe’s cost reduction is higher for Llama-3.2 (CroAttn) model as the increase in image tokens caused by image-driven bursts not overwhelming LLM backend in CroAttn models as observed in its latency profile (Figure 12). However, in InternVL (DecOnly), the LLM backend’s latency increases with the increase in image tokens due to homogeneous self-attention. Therefore, to meet SLOs, ModServe scales up the number of Text Instances for InternVL more than for Llama-3.2 during image-driven bursts (light pink in Figure 16). Overall, ModServe’s stage-aware autoscaling prevents unnecessarily scaling up LLM backends (done by vLLM due to monolith deployment) during image-driven bursts and prevents resource over-provisioning.
@@ 6. Related Work
LMM Characterization. Lee et al. (Lee et al., 2024) provides a comprehensive characterization of multimodal generation models at Meta, while we focus on multimodal inputs. Hou et al. (Hou et al., 2022) focus on traditional multimodal models employing small-scale convolutional neural networks. In contrast, our work presents a detailed analysis of multimodal input workloads on both open-source LMM models and production traces, highlighting their unique execution and workload patterns.
LMM Serving. Recent research has introduced several techniques to optimize LMM serving by addressing key inefficiencies in inference computation and memory usage. Inf-MLLM (Ning et al., 2024) employs token caching strategies and attention bias to maintain performance with long contexts while reducing KV cache memory consumption. Elastic Cache (Liu et al., 2025) utilizes an importance-driven cache merging strategy to prune KV caches efficiently during inference. Dynamic-LLaVA (Huang et al., 2024), VTW (Lin et al., 2024), and QueCC (Li et al., 2024a) present various vision token sparsification and compression techniques to dynamically reduce redundancy in vision tokens. These optimizations primarily operate at the model level, trading off computational overhead with model performance. They are orthogonal to our proposed system-level design for SLO-driven LMM serving that does not impact model performance, which can further benefit from such model-level advancements, e.g., faster image encoding through token compression.
Text-Centric LLM Serving. Recent studies have delved into disaggregating LLM prefill and decode phases for text-only LLM serving. Examples include Splitwise (Patel et al., 2024), DistServe (Zhong et al., 2024), Mooncake (Qin et al., 2024), and MemServe (Hu et al., 2024). Other optimizations for LLM serving include key-value cache management (Kwon et al., 2023), continuous batching (Yu et al., 2022), request scheduling (Qiu et al., 2024a; Patke et al., 2024; Sun et al., 2024; Agrawal et al., 2024), and energy optimization (Stojkovic et al., 2025a, b; Qiu et al., 2024b). While these optimizations can be applied in ModServe to enhance LLM backend prefill and decode efficiency, our work focuses on the unique characteristics of multimodal models.

### v3

@@ 4. ModServe Design and Implementation
Overview. Figure 13 shows ModServe’s design. A pool of Image Instances handles image preprocessing and encoding of image-text requests. The resulting image tokens are passed to a pool of Text Instances, which performs LLM prefill and decode operations to generate outputs. Text-only requests bypass the image components and are queued directly at the Text Instances. Two pools are managed by the Image and Text Pool Managers.
@@ 4.1. Offline LMM Profile Generation
When onboarding a new LMM, ModServe generates resource-performance profiles by characterizing the image encoder and LLM backend independently. This profiling runs controlled inference workloads with varying model parallelisms (e.g., TP-2 and TP-8), batch sizes, and load (i.e., image tokens per second for image encoders and prompt tokens per second for LLM backends) to capture per-stage performance characteristics. To efficiently model performance across different load conditions, ModServe profiles a set of representative load levels (up to the maximum throughput) and extrapolates the behavior for intermediate loads. The resulting profiles take load, parallelism, and batch size as inputs to predict key performance metrics, including encoding latency for Image Instances and prefill time and TBT for Text Instances. The Pool Managers use these profiles to guide model-specific or architecture-specific operational decisions (Section 4.2): (1) pool autoscaling to meet latency SLOs without overprovisioning, (2) model sharding that selects the optimal TP degree, and (3) maximum batch sizing for each stage.
@@ 4.2. Decoupled Resource Management
The Image Pool Manager maintains a pool of Image Instances, which preprocess images on CPU and encode image workloads on GPU for image-text requests. Meanwhile, the Text Pool Manager manages a pool of Text Instances responsible for the prefill and decode stages of both image-text and text-only requests. Based on model profiles (Section 4.1), each manager independently optimizes pool autoscaling, model sharding, and max batch sizing to minimize costs while meeting performance SLOs. Since connectors are extremely small in size and contribute negligible latency ( of TTFT), dedicating separate GPUs to them would underutilize resources. Instead, ModServe colocates connectors with the LLM backend, which accommodates cases where the connector consumes features from multiple encoders (Zhu et al., 2025).
Before their online operation, the initial number of Image Instances () is determined using the median image QPS multiplied by the median image encoding latency. The number of Text Instances () is set as divided by the median number of images per request, based on historical LMM inference traces. If no history is available, ModServe initially overprovisions resources to both Text/Image Instances to ensure reliability.
@@
For Image Instances, image token counts are precomputed based on the static mapping from image dimensions to the number of tokens per image (Figure 3). Autoscaling of Text Instances is based on text token load in CroAttn models but total tokens in DecOnly models due to homogeneous self-attention across modalities. Advanced autoscaling hysteresis prevention techniques (Zou et al., 2024) can be employed to avoid excessive scaling actions caused by transient workload fluctuations but are not covered in this paper.
@@
When scaling beyond a single GPU server, ModServe prioritizes autoscaling over pipeline parallelism (PP) (Shazeer et al., 2020) to maximize throughput while avoiding communication overhead, seamlessly transitioning to batch-level optimizations as needed.
@@ 4.3. Modality-Aware LMM Request Serving
Instead, ModServe routes requests by input modality. Image-text requests are assigned to Image Instances with the least image-token load. Large requests (i.e., those with more images) are consequently distributed across multiple Image Instances for parallel processing and encoding (Section 3.1), preventing degraded batching performance that would occur if all images were routed to a single instance. This effectively enables a form of request chunking (Agrawal et al., 2024a), where images in a large request can be processed in an interleaved manner with other requests, reducing HoL blocking and improving scheduling flexibility.
@@
Pool Managers continuously monitor SLO attainment and trigger pool autoscaling when the rate falls below a predefined threshold (default 0.99 with a sensitivity study in Section 5.3), ensuring adaptive resource allocation under dynamic workloads (especially in cases of unpredictable traffic). ModServe can work with state-of-the-art batch scheduling techniques (Agrawal et al., 2024a; Patel et al., 2024) to optimize TBT during the decode stage, which we leave to future work as we do not observe TBT degradation in LMM characterization (Section 3.1).
Image Token Transfer. Once image encoding is complete, ModServe transfers image tokens from Image Instances to Text Instances via a pull-based RDMA mechanism. Push-based approaches immediately transmit image tokens as they are generated, requiring premature decisions about which Text Instance should receive them (and all subsequent tokens). This design increases synchronization overhead and risks suboptimal routing, as the system operates with incomplete information about token counts and runtime load.
In contrast, our pull-based design defers transfer until all image tokens for a request are ready. This many-to-one aggregation enables the scheduler to select the target Text Instance with full information, considering factors such as queue size, prefix caching, and payloads. Once the routing decision is made, the request carries RDMA addresses of the producing Image Instances, from which the chosen Text Instance pulls the image tokens.
@@ 4.4. Implementation
We implement ModServe using 5,000 lines of Python code. We base the Text Instance on vLLM (Kwon et al., 2023) (v0.7.2), a state-of-the-art generative model inference platform, and build the Image Instance on HuggingFace Transformers (HuggingFace, 2024b). The modular architecture of ModServe enables easy integration with other serving engines (e.g., TensorRT (NVIDIA, 2024) and DeepSpeed (Aminabadi et al., 2022)). We use numactl to restrict CPU and memory usage of image preprocessing to a single NUMA node, which reduces memory access latency and performance variation. To ensure efficient GPU-to-GPU memory transfer of image tokens, we use PyTorch’s distributed communication with the NCCL backend and GPU Direct RDMA.
@@ 5.1. Experimental Setup
Models and Workloads. We use two representative multimodal models, Llama3.2-11B and InternVL-26B, for CroAttn and DecOnly LMMs, respectively. To ensure realistic workload distribution, we adopt the inter-arrival timestamps of requests and the number of images associated with each request (ranges from 0 to 16) from the production LMM inference trace (Section 3.2) and reuse the LMM text-image dataset from Section 3.1.
@@
SLO Definition. We define the SLO metrics for LMM inference based on the TTFT/TBT during the isolated run of a single text-only request and text-image request (with one image) on the monolith baseline setup. Then, we scale the SLO metrics with a constant factor (i.e., SLO factor) to evaluate how ModServe performs under tight/relaxed SLOs (Section 5.3). The SLOs are defined on P99 tail latency across requests over time.
@@ 5.2. End-to-end Performance
Figure 16 compares the number of GPUs used by ModServe and vLLM (monolith) to serve the image-driven burst in the production trace. ModServe takes 41.3% and 25% fewer GPUs compared to vLLM to serve Llama-3.2 (CroAttn) and InternVL (DecOnly) models respectively while meeting the tail latency SLOs. ModServe’s cost reduction is higher for Llama-3.2 (CroAttn) model because the increase in image tokens caused by image-driven traffic bursts does not overwhelm the LLM backends in CroAttn models as observed in its latency profile (Figure 12). However, in InternVL (DecOnly), the LLM backend’s latency increases with the increase in image tokens due to homogeneous self-attention. Therefore, to meet SLOs, ModServe scales up the number of Text Instances for InternVL more than for Llama-3.2 during image-driven bursts (light pink in Figure 16). Overall, ModServe’s stage-aware autoscaling prevents unnecessarily scaling up LLM backends (done by vLLM due to monolith deployment) during image-driven bursts and prevents resource over-provisioning.
@@ 5.3. Sensitivity Study
Model Architectures. Our evaluation on open-source LMMs includes models up to 90B parameters, while production deployments may involve even larger model sizes affecting image encoding ratios in TTFT, which we defer to future work. We focus on visual LMMs but audio-based multimodal models (HuggingFace, 2024a) share similar architectures and parameter scales with vision multimodal models. We also note that hybrid multimodal architectures have been proposed (Dai et al., 2024), though no open-source hybrid models are currently available.
5.4. Prefill-Decode Disaggregation Support
ModServe is complementary to existing techniques for LLM backend optimization, including prefill/decode (PD) disaggregation (Patel et al., 2024; Zhong et al., 2024), which splits LLM inference into two execution phases: prefill and decode (token-by-token generation). Our design fully supports PD disaggregation, which leads to a full EPD disaggregation.
To demonstrate this, we compare two deployment configurations under varying load, both incorporate PD disaggregation, deploy the InternVL-26B model, and use the same number of decode instances to match TBT latency (orthogonal to ModServe’s contributions). The main difference between the two configurations comes in the LLM prefill instances: (1) PD-Monolith: 4 prefill instances are deployed, where each instance is distributed across 8 GPUs. Each prefill instance also hosts an image encoder for image preprocessing and encoding. (2) PD-ModServe: 3 prefill instances are deployed, each across 8 GPUs. Image encoders are fully decoupled from the LLM backends and run as 8 independent processes on the remaining GPU server. Both configurations use a total of 32 GPUs for image encoding and LLM prefill combined.
This setup allows us to isolate the benefits of stage-level decoupling in ModServe from PD disaggregation. Figure 19 demonstrates that ModServe (blue) provides additional TTFT reduction (up to 2.8× in average TTFT and 3.2× in P90 TTFT for InternVL-26B) beyond what PD disaggregation alone can offer (red). The TTFT improvement (for both mean and P90) becomes more pronounced when load increases as ModServe reduces resource contention between the image encoding and LLM prefill stages and leverages encoder parallelization to reduce encoding latency (Section 3.1).
5.5. Token Transfer Overhead
Figure 20 shows the image token transfer overhead for varying-sized image embeddings, comparing different communication media of using Infiniband and Ethernet. With RDMA on Infiniband, we observed the P99 transfer latency of image tokens per image request is 5 ms, which corresponds to ¡0.5% and ¡0.3% TTFT for CroAttn and DecOnly models, respectively. TCP over Ethernet incurs significantly higher overheads, with a P50 of 100 ms and a P99 of 180 ms. ModServe supports both communication media. When evaluated over TCP, ModServe achieves a 35% TTFT reduction at high load for InternVL-26B and an 8.4% reduction at low load compared to the monolithic baselines (with Infiniband, the reduction is 46% and 13%, respectively, as mentioned in Section 5.2).
@@ 6. Related Work
LMM Characterization. Lee et al. (Lee et al., 2024) provides a comprehensive characterization of multimodal generation models at Meta, while we focus on LMMs with multimodal input (e.g., visual understanding models). Hou et al. (Hou et al., 2022) focus on traditional multimodal models employing small-scale convolutional neural networks. In contrast, our work presents a detailed analysis of multimodal input workloads on both open-source LMM models and production traces, highlighting their unique execution and workload patterns.
LMM Serving Optimization. Recent research has introduced several techniques to optimize LMM serving by addressing key inefficiencies in inference computation and memory usage. Inf-MLLM (Ning et al., 2024) employs token caching strategies and attention bias to maintain performance with long contexts while reducing KV cache memory consumption. Elastic Cache (Liu et al., 2025) utilizes an importance-driven cache merging strategy to prune KV caches efficiently during inference. Dynamic-LLaVA (Huang et al., 2024), VTW (Lin et al., 2024), and QueCC (Li et al., 2024a) present various vision token sparsification and compression techniques to dynamically reduce redundancy in vision tokens, especially for video workloads. These optimizations primarily operate at the model level, trading off computational overhead with output quality (i.e., accuracy). They are orthogonal to our proposed system-level design for inference efficiency that does not impact model accuracy, which can further benefit from such model-level advancements, e.g., faster image encoding with subsampling (Lei et al., 2021).
To optimize LMM inference, concurrent works adopt a similar stage decoupling idea (e.g., EPD (Singh et al., 2024) and HydraInfer (Dong et al., 2025)) and parallel encoding (e.g., IRP (Singh et al., 2024)). In contrast, our work extends beyond stage decoupling by incorporating stage-aware model configuration, modality-aware routing, and autoscaling, rooted in insights from a comprehensive systems analysis of production LMM inference workloads. In addition, our characterization and evaluation take a closer look at two representative LMM architectures, rather than being limited to decoder-only models.
Text-Centric LLM Serving. Recent studies have delved into disaggregating LLM prefill and decode phases for text-only LLM serving. Examples include Splitwise (Patel et al., 2024), DistServe (Zhong et al., 2024), Mooncake (Qin et al., 2024), and MemServe (Hu et al., 2024). Other optimizations for LLM serving include key-value cache management (Kwon et al., 2023), continuous batching (Yu et al., 2022), request scheduling (Qiu et al., 2024a; Patke et al., 2024; Sun et al., 2024; Agrawal et al., 2024a, b), and energy optimization (Stojkovic et al., 2025a, b; Qiu et al., 2024b). While these optimizations can be applied in ModServe to enhance LLM backend prefill and decode efficiency, our work focuses on the unique characteristics of multimodal models.
@@ 7. Conclusion
Acknowledgments
This project was partially supported in part by U.S. NSF grants NSF-2421782, NSF-2350425, NSF-2319988, NSF-2206522, Microsoft Research Faculty Fellowship 8300751, and the Commonwealth Cyber Initiative (CCI), an investment in the advancement of cyber research, innovation, and workforce development. For more information about CCI, visit cyberinitiative.org.